OpenStudy (nuccioreggie):

A set of equations is given below: Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? x = 2x + 1 x + 9 = 2x x + 1 = 2x + 9 x + 9 = 2x + 1

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

When you wanna solve for the solution, and they're both y=, you can set them equal to each others. If y=x, and x=2, then y=2, get it? So, it would be D, because it says that if y=x+9, and 2x+1=y, then x+9=2x+1